A large speaker at an outdoor concert has a diaphragm which oscillates back and forth with an amplitude of 3mm and a frequency of 100 Hertz. Air pressure waves, also known as "sound waves" are produced by this motion.

a. Are the sound waves longitudinal or transverse?

b. Are the sound waves travelling waves or standing waves?

c.From the information given, is it possible to know what the frequency of the sound waves is? If so, what is it?

d. From the information given, it is possible to know what the wavelength of the sound waves is? If so, what is it?

e.From the information given, it is possible to know what the amplitude of the sound waves is? If so, what is it?
Answer #1

sound waves are longitudinal wave , and are travelling waves

yes frequency of sound wave is equal to frequency of source = 100 Hz

and wavelength of sound can be calculated using speed of sound wave in air that is 343 m/s so wavelength = 343 /100 = 3.43 m

amplitude of sound waves is same as amplitude of diaphragm = 3mm

please rate it up thanks :)